OpenStudy (anonymous):

PLEASE HELP ME AND I WILL GIVE MEDAL AND BE SSOOOO GREATFUL <3 3.Carmen is trying to determine if the following statement is always, sometimes, or never true: The square of a prime number is odd. She writes the following statements: 32 = 9, 52 = 25, 72 = 49, 112 = 121, and 132 = 169. Using her results, she concludes that the statement is always true. a. What kind of reasoning did Carmen use? b. Is her conclusion correct? If not, use a counterexample to prove why it is not. Answer:

OpenStudy (anonymous):

@Haseeb96 Pls help me :)

OpenStudy (anonymous):

@ganeshie8 @Here_to_Help15 @study100

OpenStudy (solomonzelman):

For #3, try `2`. But on the other hand the test primes are odd, and `odd ² = odd` . For `she writes the following...` you mean \(\normalsize\color{black}{ 3^2 = 9,~~ 5^2 = 25, ~~7^2 = 49, ~~11^2 = 121, ~~\rm{and}~~ 13^2 = 169. }\) Yes, this is true :)

OpenStudy (solomonzelman):

You do the rest. I'll be offline, but many others can finish:) Good Luck !

OpenStudy (anonymous):

okay thank you. Can @Haseeb96 you please help me?

OpenStudy (anonymous):

@Here_to_Help15